The year-end 2018 balance sheet for Brads Copy, Inc. lists common stock ($1.25 par value) at $7,626,000 capital surplus at S86,543,126 and retained earnings at $218,546,280. On the 2017 year-end balance sheet, retained earnings are listed as S214,368,950. The firms net income in 2018 was $12,659,120. No stock was issued or repurchased in 2018. What were dividends per share paid by the firm in 2018? The Lunder Company has total assets of S21,542,600, current labilities of $2,547,000, and long-term liabilities of $7,410,000. The firm has 370,000 shares of common stock outstanding. Compute the firms book value per share. Referring back to the previous question, if Lunders net income equals S1,055,700 and Lunder has a P/E ratio of 12.5 (note that P/E ratio -stock price per share divided by earnings per share), what is the current price of the stock?
Answer #1

1: Total dividend = Net Income – (Closing retained earnings- opening retained earnings)

= 12659120- (218546280-214368950)

=8,481,790

Number of shares = Common stock/ Par value

= 7626000/1.25

=6100800

Dividend per share = total dividend/ Number of shares

=8,481,790/6100800

=$1.39

2: Book value per share = (Assets – liabilities)/ Number of shares

= (21542600- 2547000-7410000)/370000

=$31.31

3: EPS = Net income/ shares

=1055700/370000

=2.85

Price = P/E Ratio* EPS

= 12.5* 2.85

=$35.625